. Determine the value of the unknown in the given matrix equalities


(i)



(5 −2


−4 3 )+


(x y


w z)


=


(2 −1


3 6 )


1
Expert's answer
2022-06-09T11:08:07-0400
"\begin{pmatrix}
 5 & -2 \\
 -4 & 3
\end{pmatrix}+\begin{pmatrix}
 x & y \\
 w & z
\end{pmatrix}=\begin{pmatrix}
 2 & -1 \\
 3 & 6
\end{pmatrix}"


"\begin{pmatrix}
 5+x & -2+y \\
 -4+w &3+z
\end{pmatrix}=\begin{pmatrix}
 2 & -1 \\
 3 & 6
\end{pmatrix}"

"5+x=2=>x=-3"

"-2+y=-1=>y=1"

"-4+w=3=>w=7"

"3+z=6=>z=3"


"\begin{pmatrix}
 x & y \\
 w & z
\end{pmatrix}=\begin{pmatrix}
 -3 & 1 \\
 7 & 3
\end{pmatrix}"